OpenStudy (rahulmr):

Given pr(A)=3/5, pr(B)=1/4. Find a) Pr(A' n B) if P(A u B)=3/4 b) Pr(A' n B)if A and B are mutually exclusive. Can anyone help me out how should i do question b). Thanks.

OpenStudy (usukidoll):

Did you calculate P ( A and B) P (A U B) = P(A) +P(B) - P(A and B)

OpenStudy (rahulmr):

For a) I used 1-[A - Pr(A u B)] and i got 23/20.

OpenStudy (usukidoll):

Oh you mean this formula P(E^c) = 1 -P(E) that's for P(A^c and B^c) = 1-P(A or B)

OpenStudy (rahulmr):

hmmm..

OpenStudy (usukidoll):

I'm also studying this this semester ...

OpenStudy (rahulmr):

Are you in Year 12 as well ???

OpenStudy (usukidoll):

but we know that P(A) = 3/5 and B = 1/4 and then P( A U B) = 3/4 we could use P (A U B) = P(A) +P(B) - P(A and B) but switch the terms P(A U B) -P(A)-P(B) = P(A and B) 3/4-3/5-1/4

OpenStudy (usukidoll):

it's called university hun

OpenStudy (rahulmr):

ohhhhh

OpenStudy (usukidoll):

danggg what I was supposed to do didn't even work . According to Axiom 1 \[0 \leq P(E) \leq 1 \] the probability of an event must be between 0 and 1. Can't exceed 1 and can't have negative numbers

OpenStudy (rahulmr):

yep!

OpenStudy (rahulmr):

so how should i do question b)

OpenStudy (usukidoll):

do you know set theory?

OpenStudy (usukidoll):

(A' and B) or (A or B) that leads to (A' and A) or ( B or B) empty set or (B) which becomes B

OpenStudy (usukidoll):

we need set theory to get some equations, and then we can use Axiom 3 on it

OpenStudy (usukidoll):

because the union of events becomes the sum of events like for example P(A U B) U P(A) by Axiom 3 P(A U B) + P(A)

OpenStudy (rahulmr):

I'm not quite sure about term Axiom 3

OpenStudy (agent0smith):

You could do a venn diagram for this